OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3

OpenStudy (anonymous):

y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (calculusxy):

point-slope form: \[\large y - y_1 = m(x - x_1)\] \(m\) always represents the slope and the \(y_1\) and \(x_1\) are the places where you would need to fill it in with the x and y coordinates of \((-1,6)\).

OpenStudy (anonymous):

so it would be y-6=m(x- -1) ? sorry im not really good at math..

OpenStudy (calculusxy):

Yes you're almost there, what is the slope?

OpenStudy (anonymous):

im not sure.

OpenStudy (calculusxy):

The question explicitly says it.

OpenStudy (anonymous):

its - 3

OpenStudy (anonymous):

i know sorry i forgot

OpenStudy (calculusxy):

Good now plug that in for m.

OpenStudy (anonymous):

y-6=-3(x- -1)

OpenStudy (calculusxy):

Good, now that is fine, but if you have a minus sign and a negative, you can make it into a positive because \((- \times - = +)\). So you can just modify this and make it into \(y - 6 = -3(x + 1)\) or, if you want, you can leave it as \(y - 6 = -3(x - (-1))\) and both of them should equal to each other.

OpenStudy (anonymous):

oh yeah i forgot (- x -= +)

OpenStudy (anonymous):

Okay thank you for the help :)

OpenStudy (mathmale):

How would you now check your answer (y-6=-3(x+1) )? Take the coordinates of the given point and substitute them back into your equation (above). Is the equation then true? If so, your solution is correct. If not, it's not.

OpenStudy (danjs):

if someone wrote a master thread that always stays under the ask questions box, and showed forming linear equations , i think we may be out of business here

OpenStudy (anonymous):

Thank you all for your help

OpenStudy (calculusxy):

You're very welcome !